OpenStudy (anonymous):

A principal of $835 is invested in an account at 3 percent per quarter simple interest. Which of the following sequence(s) describes the dollar amount of principal in the account at quarterly intervals?

OpenStudy (anonymous):

\[S.I= \frac{ P \times R \times T }{ 100 }\] S.I= Simple Interest P=Principal R=Rate T=Time